OpenStudy (anonymous):

Find the value of c such that ∫_0^c e^-2x dx=1/4

geerky42 (geerky42):

First, you would want to evaluate\(\displaystyle\int_0^ce^{-2x}~\mathbb dx \). Try using u-substitution.

geerky42 (geerky42):

Let \(u = -2x\)

geerky42 (geerky42):

Can you handle it? @SARCO06

OpenStudy (anonymous):

Not actually, I'm kind of lost

OpenStudy (anonymous):

Please help me, is my last question

geerky42 (geerky42):

Alright, so since \(u = -2x\), we have \(\mathbb du = -2\mathbb dx\quad\Longrightarrow\quad -\dfrac{1}{2}\mathbb du = \mathbb dx\), right? So make substitutions and you have \(\displaystyle\int_0^c -\dfrac{1}{2}e^{u}~\mathbb du =-\dfrac{1}{2}\int_0^c e^{u}~\mathbb du \). Now can you evaluate \(\displaystyle -\dfrac{1}{2}\int_0^c e^{u}~\mathbb du \)?

geerky42 (geerky42):

Are you still lost? Where are you stuck on?

OpenStudy (anonymous):

Yes, I'm lost, I'm trying to evaluate the value of c, but I'm also working on another homework

OpenStudy (anonymous):

Are you still there?

geerky42 (geerky42):

\(\large \displaystyle \int_0^c e^u\mathbb du ~~=~~ \left.e^u\phantom{\dfrac{}{}}\right|_0^c~~=~~\left.e^{-2x}\phantom{\dfrac{}{}}\right|_0^c = (e^{-2c})-(e^0)\) Correct?

OpenStudy (anonymous):

Yes, but what's c finally?

OpenStudy (anonymous):

Sorry, but I don't get it. I have done this function with the values. I never found one of the values before

OpenStudy (anonymous):

You have been very kind and patient with me, I really need to finish this homework. I have the rest of it done

OpenStudy (loser66):